# Tropicalization through the lens of Łukasiewicz logic, with a topos theoretic perspective

### v1

@@
Tropicalization through the lens of Łukasiewicz logic, with a topos theoretic perspective
A. Di Nola, B. Gerla, G. Lenzi
@@ 1 Introduction
The main aim of this paper is to show that the topics of Łukasiewicz logic, semirings and tropical structures fruitfully meet, by combining the ideas of [10], [17], [38] and [14]. This gives rise to a topos theoretic perspective to Łukasiewicz logic, see [12]. Our aim seems to be completely in line with the spirit of the following remark from [34]:
The confluence of geometric, combinatorial and
@@
is one of the manifestations of the unity of mathematics.”
@@
Usually $S=G\cup\{\infty\}$ where $G$ is a totally ordered abelian group and $\infty$ is an infinite extra element, but we suggest to be more liberally inspired by [38]: first, $G$ can be any lattice ordered abelian group, and $S=G\cup\{\infty\}$ is what we call the Rump semifield of $G$ Moreover, we point out that the passage from $G$ to $S$ and conversely is functorial, even an equivalence. So we propose here a functorial tropicalization.
A polynomial $p$ with coefficients in $K$ is replaced by a polynomial $Trop(p)$ with coefficients in $S$. In the tropicalization $Trop(p)$ of $p$, the polynomial product of $K$ becomes $+$ in $S$, the polynomial sum in $K$ becomes the infimum in $S$. The constants $k\in K$ become their value $v(k)$. The idea is that the tropicalized polynomial $Trop(p)$ should be simpler than $p$ but retain the combinatorial structure of $p$.
@@
(or its dual with min instead of max). It is used in many contexts, from algebraic geometry to optimization to phylogenesis. An application to differential equations is given in [22]. The variety generated by $\mathbb{R}^{maxplus}$ includes the semiring reduct of $[0,1]$, but not the other way round. Instead, if we consider the negative cone of the reals, $\mathbb{R}^{neg}=(\mathbb{R}^{\leq 0}\cup\{-\infty\},max,+,-\infty,0)$, we have that the varieties generated by $[0,1]$ and $\mathbb{R}^{neg}$ are the same. Another famous semiring is the tropical semiring $\mathbb{N}^{trop}=(\mathbb{N}\cup\{+\infty\},min,+,+\infty,0)$. Once again, the varieties generated by $\mathbb{N}^{trop}$ and $[0,1]$ are the same. However, $\mathbb{N}^{trop}$ and $[0,1]$ have different first order theories. Thus, looking at these two varieties from a logical point of view, some differences emerge. This justifies approaching tropical algebraic structures from a logical point of view.
One can try to apply the same tropicalization idea to more general frameworks, like universal algebraic geometry by Plotkin [35]. This paper proposes a possible generalization of tropicalization of algebraic structures using many valued logic and especially fuzzy logic. In fact, the algebraic structures of fuzzy Łukasiewicz logic (MV-algebras) have a natural structure of idempotent semiring. Note that MV-semirings are mentioned in [20] in the context of weighted logics.
@@
Like the semantic counterpart of classical logic is given by Boolean algebras, the semantic counterpart of Łukasiewicz logic is given by MV-algebras. These are algebraic structures generalizing Boolean algebras and widely used in many applications, from quantum mechanics to games to functional analysis. The interplay between semirings and MV-algebras is very interesting. Every MV-algebra is in a natural way (better, in two natural dual ways) a semiring, as explained in [15], [14].
Namely, MV-algebras are structures $(A,\oplus,0,1,\neg)$ where $(A,\oplus,0)$ is a commutative monoid, $\neg 0=1$, $x\oplus 1=1$, $\neg\neg x=x$ and ¬($\neg(\neg x\oplus y)\oplus y)=\neg(\neg y\oplus x)\oplus x)$.
@@
The main example of MV-algebra is $[0,1]$ with $x\oplus y=min(x+y,1)$ and $\neg x=1-x$. By [13] it follows that $[0,1]$ generates the variety of MV-algebras.
[…]
@@
In order to define other interesting MV-algebras it is convienient to introduce Mundici functor $\Gamma$, see [31]. $\Gamma$ is a functorial equivalence between MV-algebras and Abelian $\ell$-groups with strong unit. This equivalence was originally used in order to clarify the relations between $AF$ $C^{*}$-algebras and their $K_{0}$ group. $\Gamma(G,u)$ is the interval $[0,u]$ of $G$ where $x\oplus y=(x+y)\wedge u$ and $\neg x=u-x$.
Another important example is the Chang MV-algebra $C=\Gamma(Z\,lex\,$ $Z,$ (1,(1,( 1 , 0))0))0 ) ) where $lex$ denotes lexicographic product of groups. We call $V(C)$ the variety generated by $C$ and $V(C)$-algebras the elements of $V(C)$. The algebra $C$ has a very special role in MV-algebras theory. $\Gamma^{-1}(C)$ is the $K_{0}$ of Behncke-Leptin $C^{*}$-algebra $A_{0,1},$ see [32].
An equivalence related to $\Gamma$ is the one called $\Delta$ between abelian $\ell$-groups and perfect MV-algebras, see ([17]), that is, MV-algebras generated by their infinitesimals. Namely, if $G$ is an abelian $\ell$-group, then $\Delta(G)=\Gamma(Z\ lex\ G,(1,0))$. So, $C$ can also be defined as $\Delta(Z)$.
We can think of the logic of perfect MV-algebras as the logic of the quasi true or quasi false.
We focus on algebraic models of an equational extension of $\L$ukasiewicz logic. Actually we consider the extension of $\L$ukasiewicz logic given by the equation in the section 3.1.
It is often stressed that enriched valuations represent ”perturbations” of valuations of classical models. In our case we can speak of ”infinitesimal perturbation” of classical propositional logic, that in turn, means infinitesimal perturbation of Boolean algebras. Although it seems rather exotic to consider infinitesimally perturbed Boolean algebras as models of a propositional logic [33], it happens that these models can have an elegant algebraic characterization. Indeed, we can see that, albeit by means of a categorical equivalence, such models can be described by a Stone space and a family of lattice ordered abelian groups. Actually we will consider weak boolean products of lattice ordered abelian groups. It will be quite interesting that such a logic may have theories interpreted into points of the presheaf topos over the multiplicative monoid of integers.
The rest of the paper is organized as follows:
In section $3$ we use the variety $V(C)$ of MV algebras generated by $C$ to give a functorial notion of tropicalization of an algebraic structure. We also present a geometric invariant $\theta$ for algebras from the variety $V(C)$ which has role in the definition of tropicalization functor.
In section $4$ we analyze the algebraic aspects of $\theta$ invariants, showing that they can be endowed with a rich algebraic structure. Such structures form a class of extended semiring (lattice-Bisemirings) and a new kind of two sorted first order structures called Coupled $\ell$-Bisemirings. The study of such structures could allow a classification of $\theta$ invariants.
In section $5$ we pass to the particular case of perfect MV-algebras, exploring the role of $\theta$ functor in moving from $V(C)$-algebraic varieties into tropical algebraic varieties.
In section $6$ we will show how a logic admitting only truth values infinitesimally close to boolean values, surprisingly, it can exhibit models that are functorially connected to a category of points of a non-commutative geometry.
In section $7$ we sketch some conclusions.
2 MV-algebras, semirings and semifields
@@
So far we mentioned the role of semirings and idempotent semifields in the tropicalization issues. Now we observe that studies concerning the relationship between semirings and MV-algebras started in [14] and developed in subsequent papers, see ([15],[18], [19], [21]). Furthermore, via semirings, in ([41]) Łukasiewicz logic is connected with Weighted Logics, introduced by Droste and Gastin ([20]). A generalisation of MV-semirings, actually the Near semirings, are used in [4] to represent a class of quantum structures.
@@
Definition 2.1.
An MV-semiring $(A,+,\odot,0,1)$ is a commutative idempotent semiring such that for every $s\in A$ there is a greatest element $s^{*}$ such that $s\odot s^{*}=0$ and we have $s+t=(s^{*}\odot(s^{*}\odot t)^{*})^{*}$ for every $s,t\in A$.
We call homomorphism of MV-semirings a homomorphism of semirings preserving *. In fact, not every homomorphism of semirings preserves *.
For instance, take the MV-algebras $\{0,1/2,1\}$ and $\{0,1\}$, take the map $h:\{0,1/2,1\}\to\{0,1\}$ such that $h(0)=0$, $h(1/2)=0$, $h(1)=1$. Then $h$ is a homomorphism of semirings, but does not preserve *, since $h((1/2^{*}))=h(1/2)=0$ and $h(1/2)^{*}=0^{*}=1$.
Definition 2.2.
A coupled semiring is a structure $(A,+,\odot,0,1,+^{\prime},\odot^{\prime},\alpha)$ such that $(A,+,\odot,0,1)$ and $(A,+^{\prime},\odot^{\prime},1,0)$ are commutative idempotent semirings, $\alpha$ is an isomorphism between them with $\alpha(\alpha(x))=x$, and $x+y=x\odot^{\prime}(\alpha(x)\odot y)$.
A morphism of coupled semirings is defined as a map which preserves all the operations of the structure. The idea of coupled semiring is extended in [9] to residuated lattices satisfying the law of double negation.
We consider the categories of MV-algebras, MV-semirings and coupled semirings, with their corresponding homomorphisms. In [2] it is shown that such categories are equivalent, we repeat here the main steps of the proof.
Theorem 2.1.
The categories of MV-algebras and MV-semirings are equivalent.
Proof.
We define a functor $S$ from MV-algebras to MV-semirings. Given an MV-algebra $A$, let $S(A)$ be the MV-semiring associated to $A$. Given a homomorphism $h:A\to B$ between two MV-algebras, then $h$ preserves $+$, $\odot$ and , so $h:S(A)\to SB$ is a homomorphism of MV-semirings and we let $Sh=h$.
Suppose $k:SA\to SB$ is a morphism of MV-semirings associated to MV-algebras $A$ and $B$. Then $k$ preserves $\odot$ and , so $k$ preserves also $\oplus$. Hence, $k$ preserves the MV-algebra structure of $A$ and $k$ is a homomorphism of MV-algebras, and $Sk=k$. Clearly if $Sk=Sk^{\prime}$ then $k=Sk=Sk^{\prime}=k^{\prime}$. So $S$ is full and faithful
Finally, every MV-semiring $(S,\odot)$ is associated to the MV-algebra $(S,0,\oplus,*)$ where $a\oplus b=(a^{*}\odot b^{*})^{*}$, so $S$ is surjective on objects.
Summing up, $S$ is an equivalence ([30]). ∎
In the same way one proves:
Corollary 2.1.
The categories of MV-algebras and coupled semirings are equivalent.
Consider now the subcategory of MV-algebras given by Boolean algebras. A Boolean algebra is an MV-algebra where $x^{2}=x$, where we use the notation $x^{2}=x\odot x$. So by restricting the above results we have:
Corollary 2.2.
The categories of Boolean algebras and MV-semirings verifying $x^{2}=x$ are equivalent. The categories of Boolean algebras and coupled semirings verifying $x^{2}=x$ are equivalent.
Now let us turn to semifields, which are our tropical objects.
A semifield is a semiring $(F,+,\times,0,1)$ where the multiplication is commutative and every nonzero element has an inverse with respect to the multiplication. A semifield is idempotent if $x+x=x$ for every $x\in F$.
@@
The equivalence in Theorem 2.2 acts by associating with every idempotent semifield $(F,+,\times,0,1)$ the abelian $\ell$-group $(F\setminus\{0\},\cdot,1)$..
The equivalence $\rho$ can be refined as follows. Let $(F,+,\times,0,1)$ be an idempotent semifield. Let $u\in F$. We say that $u$ is a strong unit of $F$ is $u\not=0$ and for every $x\in F,x\not=0$, there is $n$ such that $x^{n}\leq u$. Again, by removing the zero of the semifield we get the following:
Proposition 2.1.
There is an equivalence between idempotent semifields with a strong unit and abelian $\ell$-groups with a strong unit.
Now by composing $\Gamma$ and $\rho$ we obtain:
@@
The previous theorem supports the claim that tropical geometry can be made not only on semifields, but even on MV-algebras. That is we get a perspective for a tropical geometry of algebraic models of $\L$ukasiewicz logics.
Moreover, note that Theorem 2.3 specializes to subvarieties as follows:
Corollary 2.3.
For every subvariety $V$ of MV-algebras there is an equivalence between $V$ and a category of semifields with strong unit definable in first order logic over the language of semifields with strong unit.
Proof.
Every variety $V$ of MV-algebras is defined by a finite set $\Sigma_{V}$ of equations in one variable. Now for every abelian $\ell$-group $G$, the condition that $\Gamma(G,u)$ satisfies $\Sigma_{V}$ is definable in $G$ by a first order formula $\phi$ over the $\ell u$-group language of $G$. In turn this condition is equivalent to a first order formula $\phi^{nonz}(x)$ over $\rho^{-1}(G)$ in the language of semifields. In fact, $\phi^{nonz}(x)$ is obtained by restricting the quantifiers of $\phi$ to the nonzero elements of the semifield. ∎
By composing the functors $\Delta$ and $\rho$ we have another result concerning perfect MV-algebras (which are not a variety):
@@
Theorems 2.3 and 2.4 are the first steps of a comparison between tropical geometry and geometry over MV-algebras.
3 Rump tropicalization and its generalization via $V(C)$ algebras
Our aim is to show how much the topics of $V(C)$-algebras, their universal algebraic geometry, logic, semirings and Tropical structures fruitfully meet. We will show how a special class of semirings, which are reducts of MV-algebras, is related with issues of Dequantization (see [38]), Tropical Geometry and Non Commutative Geometry. This will be seen through the lens of Super Łukasiewicz Logic, [16], obtaining a logical and geometric view of tropical structures.
We propose a functor $\theta$ from the category $V(C)$ to $Set$ which, in our opinion, plays the role of a tropicalization of algebraic structures and extends Rump tropicalization of abelian $\ell$-groups, see [38]. We recall that Rump associates to every abelian $\ell$-group a semifield such that:
  1. 1.
    the semifield is idempotent;
  2. 2.
    the correspondence is a categorical equivalence.
We can say that the tropicalization of a category (in the sense of Rump) is a functor satisfying the two properties above.
3.1 $\theta$ invariant of $V(C)$-algebras
In this section we will focus on the equational class of $V(C)$ which is a particularly natural extension of Boolean algebras inside the variety of MV-algebras. Indeed, as has been noted in [8], the MV-algebras in $V(C)$ are Boolean algebras “perturbed” with infinitesimals. More precisely, every element is infinitely close to a Boolean element. From the point of view of universal algebra, $V(C)$ is an atom in the lattice of subvarieties of the variety of MV-algebras, where the variety $Bool$ of Boolean algebras is the smallest subvariety.
Let us define infinitesimal an element $x$ such that $nx\leq\neg x$ for every positive integer $n$. The idea is that $x\leq 1/n$, but $1/n$ does not necessarily exist in an MV-algebra. The set of all infinitesimals of an MV-algebra $A$ is called the radical of $A$ and is denoted by $Rad(A)$. A perfect MV-algebra is an MV-algebra generated by its radical, hence by its infinitesimals.
Note that $[0,1]$ has no infinitesimals, whereas $C$ has them. Indeed, $C$ consists only of infinitesimals and negations of infinitesimals, that is, is perfect. Every MV-algebra has a largest perfect subalgebra, consisting of the radical and coradical.
The co-radical of the Lindenbaum algebra of first order Łukasiewicz logic is given by the formulas which are true but not provable, see [39]. It would be interesting to describe the radical of this MV-algebra.
$(2x)^{2}=2(x^{2})$
where $2x$ means $x\oplus x$ and $x^{2}$ means $x\odot x$. We remark that the corresponding logic has some nice properties, for example, the propositional logic is complete with respect to the MV-chain $\Delta(R)$, and a first order logic is complete with respect to all Perfect MV-chains.
We note that the above identity is used in [11] in defining the variety of DL-algebras and its free algebras.
In $V(C)$, every element is infinitely close to a Boolean element, that is to an element $x=x\oplus x$. So, we can consider $V(C)$-algebras as perturbed Boolean algebras. Note that infinitesimals cannot exist in Boolean algebras, but they occur in mathematics, from nonstandard analysis [36] to first order fuzzy logic [39]. More generally, every MV-algebra $A$ has a Boolean part $B(A)$, a Boolean algebra consisting of the elements $x$ such that $x=x\oplus x$, and a perfect part $P(A)$, the perfect MV-algebra generated by the infinitesimals of $A$. If $A\in V(C)$, then $A$ is a combination of $P(A)$ and $B(A)$, a perfect MV-algebra and a Boolean algebra. But by the functor $\Delta$, a perfect MV-algebra corresponds to an Abelian $\ell$-group. So in a sense, an MV-algebra in V(C) is a combination of a Boolean algebra and an Abelian $\ell$-group. We can say that the Boolean algebra carries a qualitative information, and the $\ell$-group carries a quantitative information.

### v2

@@
Abstract
The main aim of this paper is to show the interconnections between Łukasiewicz logic and algebraic geometry using algebraic, geometric and logical instruments. We continue our investigation into a new algebraic geometry based on idempotent semifields, in particular those related with MV-algebras, describing categorical equivalence between different structures related to tropical geometry and many valued logics. Further, we describe such connections in terms of topoi.
@@ 1 Introduction
The main aim of this paper is to combine the ideas of [7], [13], [26] and [10] to show that the topics of Łukasiewicz logic, tropical structures and algebraic geometry fruitfully meet. This gives rise to a topos theoretic perspective to Łukasiewicz logic, see [8]. Our aim seems to be completely in line with the spirit of the following remark from [23]:
"The confluence of geometric, combinatorial and
@@
is one of the manifestations of the unity of mathematics."
@@
Usually $S=G\cup\{\infty\}$ where $G$ is a totally ordered abelian group and $\infty$ is an infinite extra element, but we suggest to be more liberally inspired by [26]: first, $G$ can be any lattice ordered abelian group, and $S=G\cup\{\infty\}$ is a semifield (the Rump semifield of $G$) . Moreover, we point out that the passage from $G$ to $S$ and conversely is functorial, even an equivalence. So we propose here a functorial tropicalization.
@@
(or its dual with min instead of max). The variety generated by $\mathbb{R}^{maxplus}$ includes properly the semiring reduct of $[0,1]$. On the other hand, if we consider the negative cone of the reals, $\mathbb{R}^{neg}=(\mathbb{R}^{\leq 0}\cup\{-\infty\},max,+,-\infty,0)$, we have that the varieties generated by $[0,1]$ and $\mathbb{R}^{neg}$ are the same. Another famous semiring is the tropical semiring $\mathbb{N}^{trop}=(\mathbb{N}\cup\{+\infty\},min,+,+\infty,0)$. Once again, the varieties generated by $\mathbb{N}^{trop}$ and $[0,1]$ are the same. However, $\mathbb{N}^{trop}$ and $[0,1]$ have different first order theories. Thus, looking at these two varieties from a logical point of view, some differences emerge. This justifies approaching tropical algebraic structures from a logical point of view.
One can try to apply the same tropicalization idea to more general frameworks, like universal algebraic geometry by Plotkin [24]. This paper proposes a possible generalization of tropicalization of algebraic structures using many valued logic and especially fuzzy logic. In fact, the algebraic structures of fuzzy Łukasiewicz logic (MV-algebras) have a natural structure of idempotent semiring.
@@
Like the semantic counterpart of classical logic is given by Boolean algebras, the semantic counterpart of Łukasiewicz logic is given by MV-algebras.
We focus on algebraic models of an equational extension of Łukasiewicz logic. Actually we consider the extension of Łukasiewicz logic given by the equation in the section 3.1 that generates the variety containing all perfect MV-algebras. Perfect MV-algebras are characterized by the fact that their elements are either infinitesimals or negation of infinitesimals.
[…]
Tarski Conjecture 1 Any two non-abelian free groups are elementarily equivalent. That is any two non- abelian free groups satisfy exactly the same first-order theory.
Tarski Conjecture 2 If the non-abelian free group H is a free factor in the free group G then the inclusion map is an elementary embedding, see [16].
We hence organized the rest of the paper as follows:
In section $3$ we use the variety $V(C)$ of MV algebras generated by $C$ to give a functorial notion of tropicalization of an algebraic structure. We also present a geometric invariant $\theta$ for algebras from the variety $V(C)$ which has role in the definition of tropicalization functor. We further analyze the algebraic aspects of $\theta$ invariants, showing that they can be endowed with a rich algebraic structure.
In section $4$ we will show how a logic admitting only truth values infinitesimally close to boolean values, surprisingly, it can exhibit models that are functorially connected to a category of points of a non-commutative geometry
In the last section we sketch some conclusions.
2 MV-algebras, semirings and semifields
MV-algebras are structures $(A,\oplus,0,1,\neg)$ where $(A,\oplus,0)$ is a commutative monoid, $\neg 0=1$, $x\oplus 1=1$, $\neg\neg x=x$ and ¬($\neg(\neg x\oplus y)\oplus y)=\neg(\neg y\oplus x)\oplus x)$.
@@ 2 MV-algebras, semirings and semifields
The main example of MV-algebra is $[0,1]$ with $x\oplus y=min(x+y,1)$ and $\neg x=1-x$. By [9] it follows that $[0,1]$ generates the variety of MV-algebras. Further, interpretation of formulas of Łukasiewicz logic in the MV-algebra $[0,1]$ give a complete and sound semantics for the logic.
@@
In order to define other interesting MV-algebras it is convienient to introduce Mundici functor $\Gamma$, see [21]. $\Gamma$ is a functorial equivalence between the category of MV-algebras and the category of Abelian $\ell$-groups with strong unit, i.e. an element $u\in G$, $u\geq 0$ such that for every element $g\in G$ there exists a natural number $n\geq 1$ such that $g\leq nu$.
This equivalence was originally used in order to clarify the relations between $AF$ $C^{*}$-algebras and their $K_{0}$ group. For any $\ell$-group $G$ with strong unit $u$, we get an MV-algebra $\Gamma(G,u)$ as the interval $[0,u]$ of $G$ where $x\oplus y=(x+y)\wedge u$ and $\neg x=u-x$.
In case $G$ is the lexicographic product of two copies of the additive group of integer numbers, we get Chang MV-algebra $C=\Gamma(\mathbb{Z}\,\stackrel{{\scriptstyle\mbox{\tiny{\emph{lex}}}}}{{\times}}\,$ $\mathbb{Z},$ (1,(1,( 1 , 0))0))0 ) ).
Let us define infinitesimal an element $x$ such that $nx\leq\neg x$ for every positive integer $n$. The idea is that $x\leq 1/n$, but $1/n$ does not necessarily exist in an MV-algebra. The set of all infinitesimals of an MV-algebra $A$ is called the radical of $A$ and is denoted by $Rad(A)$. A perfect MV-algebra is an MV-algebra generated by its radical, hence by its infinitesimals. So, every perfect MV-algebra $A$ is the disjoint union of its radical and its co-radical, i.e., the set of negations of elements of the radical: $A=Rad(A)\cup co-Rad(A)$.
Note that $[0,1]$ has no infinitesimals, whereas $C$ has them. Indeed, $C=\{(0,n),(1,-m)\mid n,m\in\mathbb{N}\}$ consists only of infinitesimals $(0,n)$ and negations of infinitesimals $(1,-m)$, that is, is perfect. Every MV-algebra has a largest perfect subalgebra, consisting of the radical and coradical.
$(2x)^{2}=2(x^{2})$
where $2x$ means $x\oplus x$ and $x^{2}$ means $x\odot x$. Every perfect MV-algebra belongs to $V(C)$, but in $V(C)$ there are also MV-algebras that are not perfect, for example to $C\times C$.
In general, every MV-algebra $A$ has a Boolean part $B(A)$, a Boolean algebra consisting of the elements $x$ such that $x=x\oplus x$, and a perfect part $P(A)$, the perfect MV-algebra generated by the infinitesimals of $A$. If $A\in V(C)$, then $A$ is a combination of $P(A)$ and $B(A)$, a perfect MV-algebra and a Boolean algebra. We can hence say that in $V(C)$-algebras every element is infinitely close to a Boolean element, that is to an element $x=x\oplus x$. So, we can consider $V(C)$-algebras as perturbed Boolean algebras ([12]).
Besides the categorical equivalence $\Gamma$, in [13] an equivalence $\Delta$ is established between abelian $\ell$-groups and perfect MV-algebras. Namely, if $G$ is an abelian $\ell$-group, then $\Delta(G)=\Gamma(\mathbb{Z}\ \stackrel{{\scriptstyle\mbox{\tiny{\emph{lex}}}}}{{\times}}\ G,(1,0))$. So, $C$ can also be defined as $\Delta(\mathbb{Z})$. By the functor $\Delta$, a perfect MV-algebra corresponds to an Abelian $\ell$-group. So in a sense, an MV-algebra in V(C) is a combination of a Boolean algebra and an Abelian $\ell$-group. We can say that the Boolean algebra carries a qualitative information, and the $\ell$-group carries a quantitative information.
@@
In the introduction, we mentioned the role of semirings and idempotent semifields in the tropicalization issues. Now we observe that studies concerning the relationship between semirings and MV-algebras started in [10] and developed in subsequent papers, see ([11],[14], [15], [17]).
@@
Now let us turn to semifields, which are our tropical objects. A semifield is a semiring $(F,+,\times,0,1)$ where the multiplication is commutative and every nonzero element has an inverse with respect to the multiplication. A semifield is idempotent if $x+x=x$ for every $x\in F$.
@@
The equivalence in Theorem 2.1 acts by associating with every idempotent semifield $(F,+,\times,0,1)$ the abelian $\ell$-group $(F\setminus\{0\},\cdot,1)$.
The equivalence $\rho$ can be refined as follows. Let $(F,+,\times,0,1)$ be an idempotent semifield. Let $u\in F$. We say that $u$ is a strong unit of $F$ is $u\not=0$ and for every $x\in F,x\not=0$, there is $n$ such that $x^{n}\leq u$. Again, by removing the zero of the semifield and by composing with functor $\Gamma$, we get the following:
@@
By composing the functors $\Delta$ and $\rho$ we have another result concerning perfect MV-algebras:
@@
The previous theorems support the claim that tropical geometry can be made not only on semifields, but even on MV-algebras. That is we get a perspective for a tropical geometry of algebraic models of Łukasiewicz logics.
Theorems 2.2 and 2.3 are the first steps of a comparison between tropical geometry and geometry over MV-algebras.
3 Rump tropicalization and its generalization via $V(C)$ algebras: the functor $\theta$
In [26] the author established a category equivalence between abelian $\ell$-groups and idempotent semifields. We say that the tropicalization of a category (in the sense of Rump) is an equivalence functor over idepotent semifields.
Consider now the variety $V(C)$ generated by Chang’s MV-algebra.